# The morphology and interface structure of titanium on graphene

### v1

@@ Abstract
Titanium (Ti) is an adhesion and contact metal commonly used in nanoelectronics and two-dimensional (2D) materials research. However, when Ti is deposited on graphene (Gr), we obtain dramatically different film morphology depending on the experimental conditions. Through a combination of transmission electron microscopy, Raman spectroscopy, and ab initio density functional theory calculations, we show that the most critical parameters are the number of Gr layers, the nature of the Gr support, and the deposition temperature. Particularly distinctive is the island morphology and large defect density of Ti on monolayer Gr, compared to bilayer or thicker Gr. We propose that this results from structural and mechanical differences between monolayer and thicker Gr flakes, where monolayer Gr is more flexible, exhibits larger surface roughness and therefore lower Ti diffusivity, and is more easily damaged. Our results highlight the extreme sensitivity of Ti morphology on Gr to processing and substrate conditions, allowing us to propose design rules for controlling Ti-Gr interface properties and morphology and to discuss the implications for other technologically relevant metal deposition processes.
I Introduction
@@
Figure 1: The effect of Gr layer number on the Ti-Gr interface. (a, b) TEM images of suspended (a) 1L and 2L Gr, (b) 3L Gr. Both images are from the same sample with 1 nm Ti deposited. A step edge separates the 1L and 2L regions in (a). (c) Combined diffraction patterns from regions of suspended 1L, 2L, and 3L Gr. Miller-Bravais indices are indicated. (d) DFT calculations of the atomic interface structure between Gr and the Ti(0001) surface. To highlight the interface structure, the Ti atoms at the interface are shown in green and the rest of the Ti atoms are shown in blue, with carbon atoms shown in dark brown.
@@
On two-dimensional (2D) materials, metal thin film quality is particularly critical for applications where the properties and performance of the system depend on the 2D material/metal interface. Notable examples include contact resistance to 2D materials [5, 6], proximity-induced superconductivity between topological insulators and superconductors [7], and spintronics applications [8]. On van der Waals (vdW) surfaces, the formation of a metal thin film follows steps of initial nucleation, growth of islands, and coalescence [9]. Factors that affect the diffusion distance or nucleation sites may thus impact the properties of the thin film that eventually forms. Various factors determine the metal-2D material interaction, including surface roughness that influences diffusion properties [10, 11], reactions of the metal with the substrate [12] or background gases during deposition [13], contamination from the ambient such as water vapor or hydrocarbons affecting metal morphology and diffusion properties [14], and contamination from preparation of the original surface [15, 16]. Understanding how the extrinsic factors influence metal diffusion, nucleation, and growth is necessary for fabricating well-controlled films with good structural and electronic properties. In graphene (Gr), a material with ultrahigh charge carrier mobility [17, 18], controlling thin film properties and metal-2D interactions is especially critical given the applications in high-frequency electronics [19], spintronic devices [20], wide-band photodetection [21], ballistic devices [22], and hydrogen storage [23].
@@
Of particular interest as a metal thin film is titanium (Ti), which is widely used as an adhesion layer in nanoelectronics due to the strong chemical interactions it can have with substrates, and which shows minimal interdiffusion into noble metals such as gold [24, 25]. Furthermore, Ti islands are of interest for optoelectronic applications, such as the bottom-up fabrication of plasmonic nanoparticle arrays for optical devices [26] and catalysis [27]. However, the exact nature of the interaction between Gr and Ti remains debated. For instance, literature is divided on whether or not defects form in Gr after Ti deposition [28, 29, 30]. In addition, it is possible that Ti and Gr react during deposition, forming TiC [31, 32]. However, another study found that TiC formation is due to a reaction between Ti and background gases or contamination in the vacuum system and not a reaction between the Ti and Gr [33].
@@
To elucidate the fundamental factors determining diffusion, nucleation, and growth of Ti on Gr, and to distinguish between underlying effects such as substrate, strain, and roughness, it is useful to consider metal deposition on suspended Gr, i.e., Gr not in direct contact with a substrate. This approach enables the study of intrinsic metal-Gr interactions without extraneous effects from the substrate. Suspended Gr is itself promising for applications in areas such as filtration [34], DNA sequencing [35], and catalyst support [36, 37, 38]. On suspended Gr, investigations of the early-stage growth of Ti reveal that growth takes place in the Volmer-Weber island growth mode and the initial island layers are strained to accommodate the lattice mismatch between Ti and Gr [39]
@@
Here, we examine the interaction between Ti and Gr systematically, as a function of Gr layer thickness, substrate, and temperature. This allows us to elucidate the influence of these different factors on the Ti-Gr interaction and the resulting Ti morphology and Ti-Gr interface structure. We organize our results as follows: (1) We describe the dependence of Ti island morphology and nucleation characteristics on Gr thickness, through a combination of TEM and ab initio density functional theory (DFT) calculations; (2) We determine the effect of deposition temperature on the Ti morphology; (3) We consider substrate effects by comparing results for suspended Gr and SiN-supported Gr; (4) We investigate defect formation during Ti deposition by Raman spectroscopy. (5) We summarize the effects of Gr thickness, temperature, substrate on Ti morphology and defect formation.
II Graphene thickness dependence on Ti-island morphology and nucleation characteristics
Figure 1(a, b) shows the results of depositing, by evaporation, a 1 nm thick Ti layer at room temperature onto 1L, 2L and 3L suspended Gr. Optical microscopy images of the Gr before and after transfer to the TEM sample carrier, shown in Fig. S1, allow us to correlate the location of images recorded in TEM with the thickness of the Gr at that point. We find that the Ti forms irregularly shaped islands on Gr, but in spite of their irregular appearance these islands are aligned epitaxially. An irregular island shape is expected given the low homologous temperature during these depositions [2],  0.15 $T_{\mathrm{M}}$, where $T_{\mathrm{M}}$=1941 K is the melting temperature of Ti.
The uniformity of the Ti film suggests that the growth is not dominated by a low density of heterogeneous nucleation sites such as surface contaminants. In general, to investigate the intrinsic interaction between metals and 2D materials, it is crucial that the 2D material is free from adsorbed contaminants, particularly water and polymer residues from the transfer process. The samples shown in Fig. 1 were prepared using a transfer process combined with sample annealing and metal deposition, the latter two conducted in ultra-high vacuum (UHV) conditions, to create the suspended and supported Gr samples on which we deposit Ti. This process results in the near-complete elimination [10, 40] of contamination on suspended Gr. Even so, it is important to minimize the time between Gr annealing and Ti deposition, because the Gr surface can re-contaminate even in UHV conditions [10, 41]. Further details of sample preparation and Ti deposition are described in Methods.
We now discuss the epitaxial arrangement displayed by the Ti on Gr. The Ti $\langle 2\bar{1}\bar{1}0\rangle$ directions are parallel to the Gr $\langle 10\bar{1}0\rangle$ directions as shown in the diffraction patterns in Fig. 1(c), and the Gr [1000] surface normal is parallel to the Ti [1000] surface normal.
We used a selected area aperture with a diameter of 520 nm. Therefore, each diffraction pattern include signals from several hundred Ti islands. The diffraction patterns show that each separate Ti island grows in identical alignment with the Gr. This crystal alignment between Ti and Gr is consistent with predictions from ab initio calculations that we performed (see Methods), as well as with results from literature [29, 39, 40, 42]. Figure 1(d) shows the expected atomic interface structure between Ti and Gr.
The nucleation density of Ti islands determines parameters such as density of grain boundaries and film stress and has a direct impact on the properties of the resulting thin films [1, 2, 3, 4]. For the Ti islands on suspended Gr in Fig. 1, we find that changes in nucleation density and island size are visible depending on the Gr layer number. Each area shown in Fig. 1(a, b) received the same Ti flux, yet the island size increases and nucleation density decreases with increasing Gr layer number. Figure 2(a) shows this quantitatively in a plot of the Ti-island nucleation density for 1L-3L Gr following the deposition of 1 nm Ti at room temperature, while Fig. 2(b) shows histograms of the island areas. Figure S2 and S3 show the details of the images from which we derived this data and the method used to determine the individual grain boundaries. Details of these measurements are given in Methods.
@@ II Graphene thickness dependence on Ti-island morphology and nucleation characteristics
Figure 2: Ti nucleation characteristics on suspended Gr. (a) Nucleation density measured from the sample shown in Fig. 1. (b) Histograms of Ti island areas for 1L, 2L, and 3L Gr. n is the number of measured islands. (c) Top and side view, respectively, of the most stable adsorption site, the hollow (H) site. Also indicated are the top (T) and bridge (B) sites.
@@
Table 1: Results from DFT calculations. The total, short-range, and vdW adsorption energies are for a single Ti atom on the hollow site. The interface energy is calculated for a Ti and Gr slab as outlined in Methods.
@@
where $D$ is the diffusion constant and $i$ is the critical nucleus size. We use values $N_{\mathrm{1L}}$ = 960 µm-2, $N_{\mathrm{2L}}$ = 590 µm-2, and $N_{\mathrm{3L}}$ = 520 µm-2 (Fig. 2). The critical nucleus size of metals can generally be considered small [43, 44], so we assume that $i$=1 or 2. These values lead to results $D_{1}/D_{2}$ = 0.18 and 0.33, respectively, and $D_{1}/D_{3}$ = 0.12 and 0.25, for $i$=1 and 2, respectively. Thus, the measurement shows that the diffusion coefficient of Ti on suspended Gr is 3-5 times larger for 2L Gr compared to 1L Gr, and 4-8 times larger for 3L Gr compared to 1L Gr.
@@
Such layer-dependent metal diffusion properties have previously been observed for gold on Gr [10, 45, 46, 47, 48]. The Gr layer dependence for Au diffusion was attributed to variations in surface roughness with layer number [10, 11]. The underlying mechanism involves the diffusion coefficient being reduced or increased by compressive or tensile strain, respectively [49, 50]. As a result, regions with compressive strain act as effective nucleation sites due to lower metal diffusivity in those areas. Suspended 1L Gr exhibits a root mean square roughness of about 0.1 nm [51, 52, 53], which decreases as the layer thickness increases [54]. Extending these ideas from gold to Ti, the thickness dependence on Ti island size in suspended Gr can be understood consistently in terms of thickness-dependent surface roughness.
@@
To aid in understanding the differences in the Ti nucleation density and interaction as a function of Gr layer number, we performed ab initio DFT calculations. As we show below, these calculations predict an increasing Ti-Gr interaction with increasing Gr thickness. Since experimental results in Fig. 1 and 2 demonstrate the opposite trend, that the Ti-Gr interaction decreases with increasing Gr thickness, our DFT results in fact suggest that extrinsic factors that are not considered in the DFT calculations dominate the layer-dependent Ti-Gr interaction.
@@
Our calculations first consider the single atom adhesion energy of Ti, which was calculated using $E_{\mathrm{ads}}=E_{\mathrm{(Gr/Ti)}}-E_{\mathrm{Gr}}-E_{\mathrm{Ti}}$, where $E_{\mathrm{(Gr/Ti)}}$, $E_{\mathrm{Gr}}$, and $E_{\mathrm{Ti}}$ are the energies of the 1L-3L Gr with adsorbed Ti, pristine Gr, and isolated Ti atoms with a spin configuration consistent with Hund’s rules, respectively. Given the symmetry of the Gr lattice, a Ti adatom can adopt three different adsorption configurations on the top, bridge, and hollow sites, as shown in Fig. 2(c). Our calculations show that, irrespective of the number of Gr layers, the hollow site is the most stable adsorption site for Ti with $E_{\mathrm{ads}}$ for 1L-3L given in Table 1. Adsorption energies for the other sites are given in Table S1. The values are consistent with previous calculations on 1L Gr [55] and are also consistent with adsorption trends of metals on other 2D surfaces such as MoS2 [56, 57].
@@
We next consider the interface energy between Gr and Ti. The interface energy was found by minimizing the misfit strain between Ti (0001) and Gr, with structures as shown in Fig. 1(d). The results for the interface energies are given in Table 1. As is the case for single adsorbed Ti atoms, we find that the interface energy decreases slightly with Gr thickness due to vdW interactions, and hence predicts a stronger Ti-Gr interaction with increasing Gr thickness.
Our overall conclusion from the ab initio calculations for both single Ti atoms and slabs of Ti is that the intrinsic interactions between Gr and Ti are not very sensitive to the Gr thickness, and that a slightly stronger interaction is expected when the Gr thickness increases. This is not consistent with the experimental results in Fig. 1, where weaker interaction, i.e. larger diffusivity was seen with increasing Gr layer number. Therefore, we posit that the differences in nucleation density for 1L-3L suspended Gr that are seen experimentally are due to an extrinsic factor, with surface roughness being a plausible origin.
@@
III Effect of deposition temperature on Ti morphology

### v2

@@ Abstract
Abstract: Titanium (Ti) is an adhesion and contact metal commonly used in nanoelectronics and two-dimensional (2D) materials research. However, we find that dramatically different film morphology can result when Ti is deposited on graphene (Gr), depending on the experimental conditions. Through a combination of transmission electron microscopy, atomic force micoscopy, and Raman spectroscopy, we show that the most critical parameters are the number of Gr layers, the nature of the Gr support, and the deposition temperature. Monolayer Gr is particularly distinctive in both its island morphology and its high defect density in Gr, compared to bilayer or thicker Gr. We attribute these results to the structural and mechanical differences between monolayer and thicker Gr flakes, where monolayer Gr is more flexible, exhibits larger surface roughness and therefore lower Ti diffusivity, and is more easily damaged. This is supported by ab initio density functional theory calculations which suggest that differences in the Ti interaction between monolayer and thicker Gr are due to extrinsic factors such as surface roughness. Our results highlight the extreme sensitivity of Ti morphology on Gr to processing and substrate conditions, allowing us to propose design rules for controlling Ti-Gr interface properties and morphology and to discuss the implications for other technologically relevant metal deposition processes.
Graphene, titanium, two-dimensional materials, transmission electron microscopy, nucleation, diffusion
@@
Figure 1: The effect of Gr layer number on the Ti-Gr interface. (a, b) TEM images of suspended (a) 1L and 2L Gr, (b) 3L Gr. Both images are from the same sample with 1 nm Ti deposited. A step edge separates the 1L and 2L regions in (a), indicated with a white dotted line. The white arrow indicates a dark elongated Ti island which is further described in Fig. S2. (c-e) The right halves show experimental diffraction patterns from suspended 1L, 2L, and 3L Gr, respectively. The left halves show simulated diffraction patterns from the predicted structures obtained from DFT (see also (f)). Miller-Bravais indices are indicated in (c). All scale bars are 2 nm-1. (f) DFT calculations of the atomic interface structure between 1L Gr and the Ti(0001) surface. To highlight the interface structure, the Ti atoms at the interface are shown in green and the rest of the Ti atoms are shown in blue, with carbon atoms shown in brown.
@@
On two-dimensional (2D) materials, the metal thin film quality is particularly critical for applications where the properties and performance of the system depend on the 2D material/metal interface. Notable examples include contact resistance to 2D materials [5, 6], proximity-induced superconductivity between topological insulators and superconductors [7], and spintronics applications [8]. On van der Waals (vdW) surfaces, the formation of a metal thin films follow steps of initial nucleation, growth of islands, and coalescence [9]. Factors that affect the diffusion distance or nucleation sites may thus impact the properties of the thin film that eventually forms. Various factors determine the metal-2D material interaction, including surface roughness that influences diffusion properties [10, 11], reactions of the metal with the substrate [12] or background gases during deposition [13], contamination from the ambient such as water vapor or hydrocarbons affecting metal morphology and diffusion properties [14], and contamination from preparation of the original surface [15, 16]. Understanding how the extrinsic factors influence metal diffusion, nucleation, and growth is necessary for fabricating well-controlled films with good structural and electronic properties.
In graphene (Gr), a material with ultrahigh charge carrier mobility [17, 18], controlling thin film properties and metal-2D interactions is especially critical in applications such as electronics [19, 20], spintronic devices [21], wide-band photodetection [22], ballistic devices [23], and hydrogen storage [24]. Applications involving other 2D materials may depend equally critically on the quality of metal-2D contacts [25]. In spite of its importance, forming consistent metallic contacts to Gr remains challenging [20]. Understanding the intrinsic interaction between metals and Gr is essential for improving the processes that form high-quality and homogeneous metallic films while minimizing defect formation.
@@
Of particular interest is titanium (Ti), which is widely used as an adhesion layer in nanoelectronics due to the strong chemical interactions it can have with substrates, and which shows minimal interdiffusion into noble metals such as gold [26, 27]. Furthermore, Ti islands are important for optoelectronic applications, such as the bottom-up fabrication of plasmonic nanoparticle arrays for optical devices [28] and catalysis [29]. However, the exact nature of the interaction between Gr and Ti remains debated. For instance, literature is divided on whether defects form in Gr after Ti deposition [30, 31, 32]. In addition, it is possible that Ti and Gr react during deposition, forming TiC [33, 34]. However, it has also been suggested that TiC formation is due to a reaction between Ti and background gases or contamination in the vacuum system and not a reaction between the Ti and Gr [35]. These varying reports suggest that the Ti-Gr interface properties are highly sensitive to experimental conditions, requiring systematic measurements to explore the impact of contributing factors.
@@
To elucidate the fundamental factors determining diffusion, nucleation, and growth of Ti on Gr, and to distinguish between underlying effects such as substrate, strain, and roughness, it is useful to consider metal deposition on suspended Gr, i.e., Gr not in direct contact with a substrate. This approach enables the study of intrinsic metal-Gr interactions without extraneous effects from the substrate. Suspended Gr is itself promising for applications in areas such as filtration [36], DNA sequencing [37], and catalyst support [38, 39, 40]. On suspended Gr, investigations of the early-stage growth of Ti reveal that growth takes place in the Volmer-Weber island growth mode and the initial island layers are strained to accommodate the lattice mismatch between Ti and Gr [41].
@@
Here, we examine the interaction between Ti and Gr systematically, as a function of Gr layer thickness, substrate, and temperature. This allows us to clarify the influence of these different factors on the Ti-Gr interaction and the resulting Ti morphology and Ti-Gr interface structure. We find that monolayer Gr behaves dramatically different compared to thicker Gr in two key respects: the Ti morphology and a large defect density that is present in the Gr after Ti deposition. We identify surface roughness–which in turn is affected by the choice of substrate and Gr thickness–as an important factor controlling the Ti morphology and defect formation in Gr. We present these results as follows: (1) We describe the dependence of Ti island morphology and nucleation characteristics on Gr thickness, through a combination of TEM and ab initio density functional theory (DFT) calculations; (2) We determine the effect of deposition temperature on Ti morphology; (3) We consider substrate effects by comparing results for suspended Gr and SiN-supported Gr; (4) We investigate defect formation during Ti deposition by Raman spectroscopy; (5) We summarize the effects of Gr thickness, temperature and substrate on Ti morphology and defect formation and discuss strategies aimed at forming Ti-Gr interfaces with minimized defect density and large expitaxial Ti grains.
@@
Figure 2: Ti nucleation characteristics on suspended Gr. (a) Nucleation density as measured from the images shown in Fig. S3. The total areas measured for 1L, 2L, and 3L Gr were 0.90, 0.55, 0.73 µm2, respectively. (b) Histograms of Ti island areas for 1L, 2L, and 3L Gr. n is the number of measured islands. (c) Top and side view, respectively, of the most stable adsorption site for Ti on Gr, the hollow (H) site. Also indicated are the top (T) and bridge (B) sites.
I Results/Discussion
I.1 Graphene thickness dependence on Ti-island morphology and nucleation characteristics
Figure 1(a, b) shows the results of depositing, by evaporation, a 1 nm thick Ti layer at room temperature onto 1L, 2L and 3L suspended Gr. Optical microscopy images of the Gr before and after transfer to the TEM sample carrier, shown in Fig. S1, allow us to correlate the location of images recorded in TEM with the thickness of the Gr at that point. We find that the Ti forms irregularly shaped islands on Gr, but in spite of their irregular appearance these islands are aligned epitaxially. An irregular island shape is expected given the low homologous temperature during these depositions [2], $\sim$0.15 $T_{\mathrm{M}}$, where $T_{\mathrm{M}}$=1941 K is the melting temperature of Ti. The darker, elongated Ti islands, such as that indicated by the white arrow in Fig. 1(a), appear to be locally thicker Ti (Fig. S2). Imaging at high magnification does not reveal visible damage to 1L Gr (Fig. S2); below we show different results when the deposition is performed at high temperature.
The uniformity of the Ti film suggests that the growth is not dominated by a low density of heterogeneous nucleation sites such as surface contaminants, a conclusion supported by the temperature dependence of nucleation density discussed below. Investigating the intrinsic interaction between metals and 2D materials generally requires that the 2D material is free from adsorbed contaminants, particularly water and polymer residues from the transfer process. The samples shown in Fig. 1 were prepared using a transfer process combined with sample annealing and metal deposition, the latter two conducted in ultra-high vacuum (UHV) conditions, to create the suspended and supported Gr samples on which we deposit Ti. This process results in the near-complete elimination of contamination on suspended Gr [10, 42]. Even so, it is important to minimize the time between Gr annealing and Ti deposition, because the Gr surface can re-contaminate even in UHV conditions [10, 43]. Further details of sample preparation and Ti deposition are described in Methods.
We now discuss the epitaxial arrangement displayed by the Ti on Gr. The Ti $\langle 2\bar{1}\bar{1}0\rangle$ directions are parallel to the Gr $\langle 10\bar{1}0\rangle$ directions as shown in the diffraction patterns in Fig. 1(c-e), and the Gr [1000] surface normal is parallel to the Ti [1000] surface normal.
Given the size of the selected area aperture used to acquire the diffraction patterns (diameter 520 nm), each pattern includes signals from tens or hundreds of Ti islands. These patterns show that each separate Ti island grows in identical alignment with the Gr. This crystal alignment between Ti and Gr is consistent with predictions from ab initio calculations that we performed (see Methods), as well as with results from literature [31, 41, 42, 44]. Figure 1(f) shows the expected atomic interface structure between Ti and Gr.
The nucleation density of Ti islands determines parameters such as density of grain boundaries and film stress and has a direct impact on the properties of the resulting thin films [1, 2, 3, 4]. For the Ti islands on suspended Gr in Fig. 1, we find that changes in nucleation density and island size are visible depending on the Gr layer number. Each area shown in Fig. 1(a, b) received the same Ti flux, yet the island size increases and nucleation density decreases with increasing Gr layer number. Figure 2(a) shows this quantitatively in a plot of the Ti island nucleation density for 1L-3L Gr following the deposition of 1 nm Ti at room temperature, while Fig. 2(b) shows histograms of the island areas. From this data we conclude that the difference in nucleation density between 1L Gr and 2L/3L Gr is statistically significant. Figure S3 shows the details of the images from which we derived this data and the method used to determine the individual island boundaries. Details of these measurements are given in Methods.
@@ I.1 Graphene thickness dependence on Ti-island morphology and nucleation characteristics
Table 1: Results from DFT calculations. The total, short-range, and vdW adsorption energies, listed in units of eV, are for a single Ti atom on the hollow site. The interface energy (eV/Å2) is calculated for a Ti and Gr slab as outlined in Methods.
@@
where $D$ is the diffusion constant and $i$ is the critical nucleus size. Equation 1 assumes that individual islands have not yet coalesced. Our islands are in close proximity but we can still distinguish between individual islands and determine from their shape and contrast that each island is a result of a single nucleation event. We use values $N_{\mathrm{1L}}$ = 1280 µm-2, $N_{\mathrm{2L}}$ = 590 µm-2, and $N_{\mathrm{3L}}$ = 520 µm-2 (Fig. 2). The critical nucleus size of metals can generally be considered small [45, 46], so we assume that $i$=1 or 2. These values lead to results $D_{1}/D_{2}$ = 0.07 and 0.18, respectively, and $D_{1}/D_{3}$ = 0.04 and 0.13, for $i$=1 and 2, respectively. Thus, the measurement shows that the diffusion coefficient of Ti on suspended Gr is 6-14x larger for 2L Gr compared to 1L Gr, and 8-25x larger for 3L Gr compared to 1L Gr.
@@
Such layer-dependent metal diffusion properties have previously been observed for Au on Gr [10, 47, 48, 49, 50]. The Gr layer dependence for Au diffusion was attributed to variations in surface roughness with layer number [10, 11]. The underlying mechanism involves the diffusion coefficient being reduced or increased by compressive or tensile strain, respectively [51, 52]. As a result, regions with compressive strain act as effective nucleation sites due to lower metal diffusivity in those areas. Suspended 1L Gr exhibits a root mean square roughness of about 0.1 nm [53, 54, 55], which decreases as the layer thickness increases [56]. Extending these ideas from Au to Ti, the thickness dependence on Ti island size in suspended Gr can be understood consistently in terms of thickness-dependent surface roughness.
@@
To aid in understanding the differences $N$ and interaction as a function of Gr layer number, we performed ab initio DFT calculations. As we show below, these calculations predict a slightly increasing adsorption energy for single Ti atoms on Gr, and an increasing interface energy for Ti slabs on Gr, with increasing Gr thickness. Since experimental results in Fig. 1 and 2 demonstrate the opposite trend, that the Ti-Gr interaction decreases with increasing Gr thickness, this comparison of experiment with DFT results suggests that extrinsic factors such as surface roughness that are not considered in the DFT calculations dominate the layer-dependent Ti-Gr interaction.
@@
Our calculations first consider the single atom adhesion energy of Ti, which was calculated using $E_{\mathrm{ads}}=E_{\mathrm{(Gr/Ti)}}-E_{\mathrm{Gr}}-E_{\mathrm{Ti}}$, where $E_{\mathrm{(Gr/Ti)}}$, $E_{\mathrm{Gr}}$, and $E_{\mathrm{Ti}}$ are the energies of the 1L-3L Gr with adsorbed Ti, pristine Gr, and isolated Ti atoms with a spin configuration consistent with Hund’s rules, respectively. Given the symmetry of the Gr lattice, a Ti adatom can adopt three different adsorption configurations on the top, bridge, and hollow sites, as shown in Fig. 2(c). The calculations show that, irrespective of the number of Gr layers, the hollow site is the most stable adsorption site for Ti with $E_{\mathrm{ads}}$ for 1L-3L given in Table 1. Adsorption energies for the other sites are given in Table S1. The values are consistent with previous calculations on 1L Gr [57] and are also consistent with adsorption trends of metals on other 2D surfaces such as MoS2 [58, 59].
@@
We next consider the interface energy between Gr and Ti. The interface energy was found by minimizing the misfit strain between Ti (0001) and Gr, with structures as shown in Fig. 1(f). Simulated diffraction patterns from such structures match experimental diffraction patterns well (Fig. 1(c-e)). The results for the interface energies are given in Table 1. As is the case for single adsorbed Ti atoms, we find that the interface energy decreases slightly with Gr thickness due to vdW interactions, and hence predicts a stronger Ti-Gr interaction with increasing Gr thickness.
Our overall conclusion from the ab initio calculations for both single Ti atoms and slabs of Ti is that the intrinsic interactions between Gr and Ti are not very sensitive to the Gr thickness, and that a slightly stronger interaction is expected when the Gr thickness increases. Our experimental observation that $N$ decreases with increasing layer thickness (Fig. 1) implies a larger Ti diffusivity on thicker Gr. This might suggest a weaker Ti-Gr interaction with increasing thickness, but this was not reproduced by DFT calculations. We therefore conclude that the experimental differences in $N$ for 1L-3L suspended Gr are due to an extrinsic factor, with surface roughness being a plausible origin.
@@
I.2 Effect of deposition temperature on Ti morphology